40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3202     Entry time: Tue Jul 13 10:02:30 2010
Author: steve 
Type: Update 
Category: MOPA 
Subject: laser power is dropping slowly 

I have just removed an other 400 cc of water from the chiller.  I have been doing this since the HTEMP started fluctuating.

The Neslab bath temp is 20.7C, control room temp 71F

 

Attachment 1: power100d.jpg  96 kB  | Show | Show all
ELOG V3.1.3-